FACTORISE- a 2 x 2 +(ax 2 +1)x+a

FACTORISE-
a2x2+(ax2+1)x+a

Grade:9

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given a2x2+(ax2+1)x+a
=a2x2+ax3+x+a
=ax3+a2x2+x+a
=ax2(x+a)+1(x+a)
=(ax2+1)(x+a)
Hence (ax2+1) and (x+a) are the factors for the above given expression